a. Using the data below as of 12/31/17 and the format found in Exhibit 4.1, create a Balance Sheet for XYZ Hospital for 12/31/17
Accounts payable 401,760
Accrued expenses 485,520
Cash 968,160
Inventory 295,600
Long-term debt 6,825,760
Long-term investments 3,844,720
Net patient accounts receivable 2,280,720
Net property and equipment 4,196,000
Notes payable 346,720
Other long-term liabilities 420,000
Short-term investments

800,000

Based upon your Balance Sheet, answer the following in the space provided:
b. How much does the hospital own?
c. What % of the hospital's assets are current?
d. How much does the hospital owe to others?
e. What are current assets minus current liabilities?
f. What is the hospital's net assets (equity)?
g. If the hospital earned $250,000 in 2017, what was the net assets (equity) balance on 12/31/16?
h. If the net assets (equity) balance was $375,000 on 12/31/16, what was the hospital's net income in 2017?
Answer #1

a). Preparing balance Sheet for XYZ Hospital for 12/31/17

Particular Amount
ASSETS
Cash 968160
Inventory 295600
Net patients Accounts Receivables 2280720
Short Term Investments 800000
Total Current Assets 4344480
Long Term Investments 3844720
Net Property & Equipment 4196000
Total Assets 12,385,200
LIABILITIES & EQUITIES
Accounts Payable 401760
Accrued Expenses 485520
Notes Payable 346720
Total Current Liabilities 1234000
Long Term Liabilities 6825760
Other Long term liabilities 420000
Total Equity (balancing figure) 3905440
Total Liabilities & Equities 12,385,200

b). Total Amount the Hospital Own = Total Assets

= 3905,440

So, Hospital own amount is $ 3905,400

c).  % of the hospital's assets that are current = (Total Current Assets / Total Assets)*100

= (4344,480/12385,200)*100

= 35.08%

So, % of hospital's assets that are current is 35.08.

d). Total Amount Hospital owe's to others = Total Current liabilities + Long Term Liabilities + Other Long term liabilities

  = 1234,000 + 6825,760 + 420,000

= $ 8479,760

So, Hospital owe amount is $ 8479,760.

e). Current Assets less Current Liabilitiies = 4344,480 - 1234,000

=$ 3110,480

f). Hospital's net assets (equity) = Total Assets - Total Liabilities

= 12385,200 - (1234,000 + 6825,760 + 420,000)

  = $ 3905,440

g). If the hospital earned $250,000 in 2017, the net assets (equity) balance on 12/31/16 is :

= 3905,440 - 250000

= $ 3655,440

h). If the net assets (equity) balance was $375,000 on 12/31/16, the hospital's net income in 2017 is :

= 3905,440 - 375,000

= $ 3530,440

Note- If you need any clarification regarding this solution, then you can ask in comments

If you like my answer then please Up-vote as it will be motivating.